### v1

# A Chung-Fuchs type theorem for skew product dynamical systems

Abstract.
We prove a Chung-Fuchs type theorem for skew product dynamical systems such that for a measurable function on such a system, if its Birkhoff average converges to zero almost surely, and on typical fibres its Birkhoff sums have a non-trivial independent random walk component, then its associated generalised random walk oscillates, that is the supreme of the random walk equals to +$\infty$+ ∞ and the infimum equals to $\infty$- ∞.
1. Introduction
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space. Let $\{W_{n}\}_{n\geq 1}$ be a sequence of independent and identically distributed (i.i.d.) random variables in $\mathbb{R}$. Let
$S=\{S_{n}=W_{1}+\cdots+W_{n}\}_{n\geq 1}$
denote its associated random walk. We say that $S$ is recurrent if for all $c>0$,
$\mathbb{P}(|S_{n}|\leq c\text{ for infinitely many }n)=1.$
We have the following well-known theorem of Chung and Fuchs [4]:
Theorem 1.1 (Chung and Fuchs 1951).
If $\mathbb{E}(|W_{1}|)<\infty$ and $\mathbb{E}(W_{1})=0$, then $S$ is recurrent. Furthermore, if $\mathbb{P}(W_{1}\neq 0)>0$, then $S$ is oscillating, that is
(1.1) $\mathbb{P}(\liminf_{n\to\infty}S_{n}=-\infty)=\mathbb{P}(\limsup_{n\to\infty}S_{n}=+\infty)=1.$
In [4], the condition “$\mathbb{E}(|W_{1}|)<\infty$ and $\mathbb{E}(W_{1})=0$” is to ensure that one can apply the strong law of large numbers to obtain
(1.2) $\mathbb{P}(\lim_{n\to\infty}\frac{S_{n}}{n}=0)=1.$
So Chung and Fuchs’ theorem can be stated in a slightly more general way without $L^{1}$-integrable assumption that
(1.2)$\ \Rightarrow\ S\text{ is recurrent}.$
For oscillation (1.1), Chung and Fuchs showed that the set of recurrent points is a closed additive subgroup of $\mathbb{R}$, hence it is either empty, or the whole space $\mathbb{R}$, or a lattice $a\cdot\mathbb{Z}$ for some $a\in\mathbb{R}$. So if $S$ is recurrent, then the set of recurrent points is non-empty, thus it is either $\mathbb{R}$ or $a\cdot\mathbb{Z}$ for some $a\neq 0$ (since $\mathbb{P}(W_{1}\neq 0)>0$), hence either case implies (1.1).
One may consider the Chung-Fuchs theorem for more general random walks: Let $(X,\mathcal{B},T,\mu)$ be a measure-preserving dynamical system (m.p.d.s) with $\mu(X)=1$. Let $f:X\to\mathbb{R}$ be a measurable function. Denote by
$S_{f}=\{S_{n}f=f+f\circ T+\cdots+f\circ T^{n-1}\}_{n\geq 1}$
its associated sequence of Birkhoff sums, or in other words, its associated generalised random walk. We say that $S_{f}$ is recurrent if for all $c>0$,
$\mu(|S_{n}f|\leq c\text{ for infinitely many }n)=1.$
The following extension of Chung-Fuchs theorem to generalised random walks is essentially due to Dekking [5] (see also [2] when assuming $f\in L^{1}(\mu)$). For readers’ convenience we shall present a proof in the next section.
Theorem 1.2 (Dekking 1982).
If $\mu$ is ergodic and
(1.3) $\mu(\lim_{n\to\infty}\frac{S_{n}f}{n}=0)=1,$
then $S_{f}$ is recurrent.
For generalised random walks, the statement that (1.3) plus $f\not\equiv 0$ implies (1.1) is no longer true. A simple counter example is the following: Let $X=\{x_{0},x_{1}\}$ be a space of two points and let the transformation $T$ be $T(x_{0})=x_{1}$, $T(x_{1})=x_{0}$ and let $\mu(\{x_{0}\})=\mu(\{x_{1}\})=\frac{1}{2}$. Take $f:X\to\mathbb{R}$ to be $f(x_{0})=1$, $f(x_{1})=-1$. Then $S_{n}f$ is periodic and (1.3) holds, but both liminf and limsup of $S_{n}f$ are bounded.
It is natural to ask for which settings between random walks and generalised random walks it holds that (1.3) plus $f\not\equiv 0$ implies (1.1). In [3] the authors studied the action of Mandelbrot cascades on ergodic measures, and in the estimation of the degeneracy of this action it was sufficient to show that
(1.4) $\mathbb{Q}(\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty)=1,$
where, in brief, the probability measure $\mathbb{Q}$ is a coupling of $\mathbb{P}$ and $\mu$, $S_{n}f(\cdot)$ is a generalised random walk under $\mu$ and $S_{n}F(x,\cdot)$ is a random walk under $\mathbb{Q}_{x}$ for $\mu$-a.e. $x$, where $\mathbb{Q}_{x}$ is the disintegration of $\mathbb{Q}$ w.r.t. $\mu$. In the sub-critical case when $S_{n}f+S_{n}F$ has a positive drift, that is $\mathbb{Q}$-a.s. $\lim_{n\to\infty}\frac{S_{n}f+S_{n}F}{n}=c>0$, it is easy to deduce (1.4). The critical case when $\mathbb{Q}(\lim_{n\to\infty}\frac{S_{n}f+S_{n}F}{n}=0)=1$ corresponds to (1.3) and it becomes more delicate to verify (1.4), which has led us to derive the following result (a slightly more general and more abstract version of the result in [3]).
Main result
Consider the skew-product measure-preserving dynamical system
$(X\times\Omega,\hat{\mathcal{B}},\hat{T},\mathbb{Q}),$
where
    ^$\hat{\mathcal{B}}=\mathcal{B}\otimes\mathcal{A}$.
    $\hat{T}(x,\omega)=(T(x),g_{x}(\omega))$ is a skew-product transformation, where for $x\in X$, $g_{x}:\Omega\to\Omega$ is $\mathcal{A}$-measurable and for $A\in\mathcal{A}$, $x\to g_{x}(A)$ is $\mathcal{B}$-measurable.
    $\mathbb{Q}$ is a coupling of $\mu$ and $\mathbb{P}$, that is $\Pi_{X}^{*}(\mathbb{Q})=\mu$ and $\Pi_{\Omega}^{*}(\mathbb{Q})=\mathbb{P}$, where $\Pi_{X}$ and $\Pi_{\Omega}$ are the projection from $X\times\Omega$ onto $X$ and $\Omega$ respectively.
Consider the measurable partition
$\eta=\{\Pi_{X}^{-1}(x)=\{x\}\times\Omega:x\in X\}$
and let $\mathbb{Q}^{\eta}_{x,\omega}$ denote the conditional measure of $\mathbb{Q}$ w.r.t. $\eta$. For $\mathbb{Q}$-a.e. $(x,\omega)$, $\mathbb{Q}^{\eta}_{x,\omega}$ is a probability measure carried by $\{x\}\times\Omega$. Note that $\mathbb{Q}^{\eta}_{x,\omega}$ only depends on $x$. Hence, equivalently, we may define a family of probability measures $\mathbb{Q}_{x}=\Pi_{\Omega}^{*}(\mathbb{Q}^{\eta}_{x,\omega})$ on $\Omega$ for $\Pi_{X}^{*}(\mathbb{Q})=\mu$-a.e. $x\in X$ such that for any measurable function $F:X\times\Omega\to\mathbb{R}$,
$\int_{X\times\Omega}F(x,\omega)\,\mathbb{Q}(\mathrm{d}(x,\omega))=\int_{X}\int_{\Omega}F(x,\omega)\,\mathbb{Q}_{x}(\mathrm{d}\omega)\mu(\mathrm{d}x).$
Let $F:X\times\Omega\to\mathbb{R}$ be a measurable function such that
  • (A)
    for $\mu$-a.e. $x\in X$, the sequence
    $\{F(x,\cdot),F\circ\hat{T}(x,\cdot),F\circ\hat{T}^{2}(x,\cdot),\cdots\}$
    is an i.i.d. sequence of random variables under $\mathbb{Q}_{x}$ such that there exist two disjoint intervals $I_{x},J_{x}\subset(0,\infty)$ such that $\mathbb{Q}_{x}(F(x,\cdot)\in I_{x})\cdot\mathbb{Q}_{x}(F(x,\cdot)\in J_{x})>0$.
For $n\geq 1$ denote by
$S_{n}F=F+F\circ\hat{T}+\cdots+F\circ\hat{T}^{n-1}.$
Let $f:X\to\mathbb{R}$ be a measurable function on $X$. One may view $f$ as a measurable function on $X\times\Omega$ by setting $f(x,\omega)=f(x)$. We have the following extension of Chung and Fuchs theorem.
Theorem 1.3.
Assume (A). If $\mu$ is ergodic and
$\mu(\lim_{n\to\infty}\frac{S_{n}f}{n}=0)=\mathbb{Q}(\lim_{n\to\infty}\frac{S_{n}F}{n}=0)=1,$
then
$\mathbb{Q}(\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty)>0.$
Remark 1.4.
If we further assume that $\mathbb{Q}$ is $\hat{T}$-ergodic, then, since $\{\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty\}$ is a $\hat{T}$-invariant set, Theorem 1.3 implies $\mathbb{Q}({\displaystyle\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty})=1$. If we also assume $-F$ satisfies (A) then $\mathbb{Q}({\displaystyle\liminf_{n\to\infty}S_{n}f+S_{n}F=-\infty})=1$.
Remark 1.5.
Assumption (A) holds for a more general class of coupling measures $\mathbb{Q}$ than the Peyrière measure studied in [3]. In particular we are not requiring that the law of $F(x,\cdot)$ under $\mathbb{Q}_{x}$ to be independent of $x$.
Remark 1.6.
In the proofs we added more details comparing to [3]. First we provided a more detailed proof of Dekking’s recurrence theorem for our need. Then we proved Lemma 2.1 to ensure that under the recurrence assumption a “suspension flow” induced from an infinite dynamical system is still a m.p.d.s.. We also added an induced system argument to deduce that a measurable (not necessarily integrable) coboundary $g=f-f\circ T$ with $f,g\geq 0$ is equal to zero almost everywhere.
2. Proofs
2.1. Proof of Theorem 1.2
Proof.
For $c>0$ and $x\in X$ define
$\tau_{c}^{0}(x)=\inf\{n\geq 1:|S_{n}f(x)|\leq c\},$
and for $k\geq 1$ define
$\tau_{c}^{k}(x)=\inf\{n\geq\tau_{c}^{k-1}(x)+1:|S_{n}f(x)|\leq c\}.$
It is equivalent to prove that for all $k\geq 0$ and $c>0$, $\tau_{c}^{k}(x)<\infty$ for $\mu$-a.e. $x\in X$.
Set $\tau^{-1}_{c}(x)=0$. Hence for $k=0$, $\tau^{k-1}_{c}(x)<\infty$ for all $c>0$ and $\mu$-a.e. $x\in X$.
We shall prove the statement by induction. Assume that for all $c>0$, $\tau^{k-1}_{c}(x)<\infty$ for $\mu$-a.e. $x\in X$. Fix $a,b,c>0$ such that $a+b=c$. For $n\geq 1$ denote by
$R_{n}(x)=\lambda\Big{(}\bigcup_{l=1}^{n}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x))\Big{)},$
where $\lambda$ is the Lebesgue measure on $\mathbb{R}$ and $I_{a}(u)=[u-a/2,u+a/2]$ stands for the interval of length $a$ centered at $u\in\mathbb{R}$. By the transition invariance of Lebesgue measure we have that $R_{n}$ is a subadditive function:
$R_{n+m}(x)\leq$ $\lambda\Big{(}\bigcup_{l=1}^{n}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x))\Big{)}+\lambda\Big{(}\bigcup_{l=n+1}^{n+m}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x))\Big{)}$
$=$ $\lambda\Big{(}\bigcup_{l=1}^{n}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x))\Big{)}+\lambda\Big{(}\bigcup_{l=n+1}^{n+m}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(x)-S_{\tau^{k-1}_{b}(x)+n}f(x))\Big{)}$
$=$ $R_{n}(x)+\lambda\Big{(}\bigcup_{l=1}^{m}I_{a}(S_{\tau^{k-1}_{b}(x)+l}f(T^{n}(x)))\Big{)}$
$=$ $R_{n}(x)+R_{m}(T^{n}x).$
Thus by Kingman’s subadditive ergodic theorem we have that for $\mu$-almost every $x\in X$,
$\lim_{n\to\infty}\frac{R_{n}(x)}{n}=\inf_{n\geq 1}\frac{\mathbb{E}_{\mu}(R_{n})}{n}.$
Using again the transition invariance of Lebesgue measure we get that
$R_{n+1}(x)=$ $\lambda\Big{(}\bigcup_{l=1}^{n+1}I_{a}(S_{\tau^{k-1}_{b}f(x)+l}f(x))\Big{)}$
$=$ $\lambda\Big{(}I_{a}(S_{\tau_{b}^{k-1}(x)+1}(x))\cup\bigcup_{l=1}^{n}I_{a}(S_{\tau_{b}^{k-1}(x)+1}f(x)+S_{l}f(T^{\tau_{b}^{k-1}(x)+1}(x)))\Big{)}$
$=$ $\lambda\Big{(}I_{a}(0)\cup\bigcup_{l=1}^{n}I_{a}(S_{l}f(T^{\tau^{k-1}_{b}(x)+1}(x)))\Big{)}$
$=$ $\lambda\Big{(}I_{a}(0)\cup\bigcup_{l=1}^{n}I_{a}(S_{l}f(T_{b,k}(T(x))))\Big{)},$
where we denote by
$T_{b,k}(x)=T^{\tau^{k-1}_{b}(x)}(x).$
By the $T$-invariance of $\mu$ and Fubini’s theorem we deduce that
$\mathbb{E}_{\mu}(R_{n+1})-\mathbb{E}_{\mu}(R_{n})$
$=$ $\int_{X}\lambda\Big{(}I_{a}(0)\cup\bigcup_{l=1}^{n}I_{a}(S_{l}f(T_{b,k}(x)))\Big{)}-\lambda\Big{(}\bigcup_{l=1}^{n}I_{a}(S_{l}f(T_{b,k}(x)))\Big{)}\,\mu(\mathrm{d}x)$
$=$ $\int_{X}\lambda\Big{(}I_{a}(0)\setminus\bigcup_{l=1}^{n}I_{a}(S_{l}f(T_{b,k}(x)))\Big{)}\,\mu(\mathrm{d}x)$
$=$ $\int_{X}\int_{I_{a}(0)}\prod_{l=1}^{n}\mathbf{1}_{\{y\not\in I_{a}(S_{l}f(T_{b,k}(x)))\}}\,\lambda(\mathrm{d}y)\,\mu(\mathrm{d}x)$
$=$ $\int_{X}\int_{I_{a}(0)}\prod_{l=1}^{n}\mathbf{1}_{\{S_{l}f(T_{b,k}(x))\not\in I_{a}(y)\}}\,\lambda(\mathrm{d}y)\,\mu(\mathrm{d}x)$
$=$ $\int_{I_{a}(0)}\mu\big{(}x:S_{1}f(T_{b,k}(x))\not\in I_{a}(y),\cdots,S_{n}f(T_{b,k}(x))\not\in I_{a}(y)\big{)}\,\lambda(\mathrm{d}y).$
This means that $\mathbb{E}_{\mu}(R_{n+1})-\mathbb{E}_{\mu}(R_{n})$ decreasingly converges to
$\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}(x))\not\in I_{a}(y)\text{ for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y)$
as $n\to\infty$. Therefore
$\inf_{n\geq 1}\frac{\mathbb{E}_{\mu}(R_{n})}{n}=$ $\inf_{n\geq 1}\frac{\sum_{m=0}^{n-1}\mathbb{E}_{\mu}(R_{m+1})-\mathbb{E}_{\mu}(R_{m})}{n}$
$=$ $\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}(x))\not\in I(y)\text{ for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y).$
Thus for $\mu$-almost every $x\in X$,
$\lim_{n\to\infty}\frac{R_{n}(x)}{n}=\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}(x))\not\in I_{a}(y)\text{ for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y).$
By the assumption that $\lim_{n\to\infty}\frac{1}{n}S_{n}f(x)=0$ and $\tau^{k-1}_{b}(x)<\infty$ for $\mu$-a.e. $x\in X$, we get that for $\mu$-almost every $x\in X$,
$\lim_{n\to\infty}\frac{S_{\tau^{k-1}_{b}(x)+n}f(x)}{n}=$ $\lim_{n\to\infty}\frac{S_{\tau^{k-1}_{b}(x)+n}f(x)}{\tau^{k-1}_{b}(x)+n}$
$=$ $0$
Thus for any $\epsilon>0$ there exists $n_{\epsilon}(x)$ such that for all $n\geq n_{\epsilon}(x)$, $S_{\tau^{k-1}_{b}(x)+n}f(x)\in[-\epsilon n,\epsilon n]$. This means that for $n\geq n_{\epsilon}(x)$,
$R_{n}(x)\leq R_{n_{\epsilon}(x)}(x)+\lambda([-\epsilon n-a,\epsilon n+a])=R_{n_{\epsilon}(x)}(x)+2\epsilon n+2a.$
Hence
$\lim_{n\to\infty}\frac{R_{n}(x)}{n}\leq 2\epsilon.$
Since $\epsilon$ is arbitrary, we get that for $\mu$-almost every $x\in X$,
$\lim_{n\to\infty}\frac{R_{n}(x)}{n}=0.$
This implies that for $\mu$-almost every $x\in X$,
$\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}(x))\not\in I_{a}(y)\text{ for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y)=0.$
Note that for $y\in I_{a}(0)=[-a/2,a/2]$ we have $I_{a}(y)=[y-a/2,y+a/2]\subset[-a,a]$. Furthermore, by the definition of $\tau_{b}^{k-1}(x)$ we have $S_{\tau_{b}^{k-1}(x)}f(x)\in[-b,b]$. Then by
$S_{\tau^{k-1}_{b}(x)+l}f(x)=S_{\tau^{k-1}_{b}(x)}f(x)+S_{l}f(T_{b,k}x),$
where when $k=0$ we may use the convention that $S_{0}f(x)=0$, we deduce that
$0=$ $\int_{I_{a}(0)}\mu\big{(}x:S_{l}f(T_{b,k}x)\not\in I_{a}(y)\text{for }l=1,2,\cdots\big{)}\,\lambda(\mathrm{d}y)$
$\geq$ $\lambda(I_{a}(0))\cdot\mu\big{(}x:S_{l}f(T_{b,k}x)\not\in[-a,a]\text{ for all }l=1,2,\cdots\big{)}$
$\geq$ $a\cdot\mu\big{(}x:S_{n}f(x)\not\in[-a-b,a+b]\text{ for all }n\geq\tau^{k-1}_{b}(x)+1\big{)}$
$=$ $a\cdot\mu\big{(}x:S_{n}f(x)\not\in[-c,c]\text{ for all }n\geq\tau^{k-1}_{b}(x)+1\big{)}$
$\geq$ $a\cdot\mu\big{(}x:S_{n}f(x)\not\in[-c,c]\text{ for all }n\geq\tau^{k-1}_{c}(x)+1\big{)}.$
This implies that $\mu\big{(}x:S_{n}f(x)\not\in[-c,c]\text{ for }n\geq\tau^{k-1}_{c}(x)+1\big{)}=0$, or in other words, for $\mu$-almost every $x\in X$,
$\tau_{c}^{k}(x)=\inf\{n\geq\tau^{k-1}_{c}(x)+1:S_{n}f(x)\in[-c,c]\}<\infty.$
2.2. Proof of Theorem 1.3
Consider the transformation $T_{f}$ on $X\times\mathbb{R}$:
$T_{f}(x,t)=(Tx,t+f(x)).$
Recall that $\lambda$ denotes the Lebesgue measure on $\mathbb{R}$. It is easy to verify that the product measure $\theta=\mu\times\lambda$ is $T_{f}$-invariant, thus $(X\times\mathbb{R},\mathcal{B}\otimes\mathcal{B}(\mathbb{R}),T_{f},\theta)$ forms a measure-preserving dynamical system. But the measure $\theta$ is a $\sigma$-finite measure with infinite mean, hence we need to induce this system to a finite one. In order to do so we need the system to be conservative. By Theorem 1.2 this is granted by (1.3).
Lemma 2.1.
If (1.3) holds and $\mu$ is ergodic then $T_{f}$ is conservative.
Proof.
It is enough to show that for set $B\in\mathcal{B}$ with $\mu(B)>0$ and $(a,b)\subset\mathbb{R}$, for every $n\geq 1$ there exists $m\geq n$ such that, writing $A=B\times(a,b)$,
$\theta(A\cap T_{f}^{-m}A)>0.$
Fix $B\in\mathcal{B}$ with $\mu(B)>0$. The induced system $(B,\mathcal{B}|_{B},T^{B},\mu_{B})$ is measure-preserving and ergodic (see [7, Proposition 1.4.1] for example), where $\mu_{B}=\frac{1}{\mu(B)}\mu|_{B}$ and $T^{B}(x)=T^{\rho_{B}(x)}(x)$ for $x\in B$, where
$\rho_{B}(x)=\inf\{n\geq 1:T^{n}(x)\in B\}$
is the first return time of $x\in B$. By Kac formula (see [7, Theorem 1.2.2] for example) we have that $\rho_{B}$ is integrable and $\int_{B}\rho_{B}\,\mathrm{d}\mu\leq 1$. Write $S_{n}^{B}\rho_{B}(x)=\rho_{B}(x)+\cdots+\rho_{B}\circ(T^{B})^{n-1}(x)$. Then Birkhoff ergodic theorem implies that
(2.1) $\lim_{n\to\infty}\frac{S_{n}^{B}\rho_{B}(x)}{n}=\tau(B):=\int_{X}\rho_{B}(x)\,\mu_{B}(\mathrm{d}x)$
exists for $\mu_{B}$-a.e. $x\in B$ and $\tau(B)\leq\frac{1}{\mu(B)}<\infty$.
For $x\in B$ write $f^{\rho_{B}}(x)=f(x)+\cdots+f\circ T^{\rho_{B}(x)-1}(x)$ and
$S_{n}^{B}f^{\rho_{B}}(x)=f^{\rho_{B}}(x)+\cdots+f^{\rho_{B}}\circ(T^{B})^{n-1}(x).$
Then (1.3) and (2.1) implies that for $\mu_{B}$-a.e. $x\in B$,
$\lim_{n\to\infty}\frac{S_{n}^{B}f^{\rho_{B}}(x)}{n}=\lim_{n\to\infty}\frac{S_{n}^{B}f^{\rho_{B}}(x)}{S_{n}^{B}\rho_{B}(x)}\frac{S_{n}^{B}\rho_{B}(x)}{n}=0.$
In other words, the measurable function $f^{\rho_{B}}$ satisfies (1.3) in terms of the induced system $(B,\mathcal{B}|_{B},T^{B},\mu_{B})$. Hence by Theorem 1.2 the generalised random walk $\{S_{n}^{B}f^{\rho_{B}}\}_{n\geq 1}$ is recurrent in this induced system.
Take $c>0$ such that $a+c<b-c$. For $(x,t)\in B\times(a+c,b-c)$ and $n\geq 1$, by recurrence for $\mu_{B}$-a.e. $x\in B$ one can find $k(x)\geq 1$ such that $S_{k(x)}^{B}\rho_{B}(x)\geq n$ and $S_{k(x)}^{B}f^{\rho_{B}}(x)\in(-c,c)$. Since $\mu(B)>0$, there exists $m\geq n+1$ such that $\mu_{B}(S_{k(x)}^{B}\rho_{B}(x)=m)>0$. Therefore
$\theta(A\cap T_{f}^{-m}A)$ $\,\geq\theta((x,t)\in B\times(a+c,b-c):S_{k(x)}^{B}\rho_{B}(x)=m)$
$\,\geq\mu(x\in B:S_{k(x)}^{B}\rho_{B}(x)=m)\cdot(b-a-2c)>0.$
Now fix $E=X\times[-c,c]$ for some $c>0$. By conservativeness (Lemma 2.1) the induced system $(E,\mathcal{B}_{E},T_{f}^{E},\theta_{E})$ is a m.p.d.s. with $\theta_{E}(E)=1$ (see [1, Proposition 1.5.3] for example). The measure $\theta_{E}$ takes the form
$\theta_{E}=\frac{1}{\theta(E)}\theta\times\lambda|_{E}=\mu\times\lambda_{c},$
where we denote by $\lambda_{c}=\frac{1}{2c}\lambda|_{[-c,c]}$ the normalised Lebesgue measure on $[-c,c]$. We may now define the following induced skew-product dynamical system
$(\overline{X},\overline{\mathcal{B}},\overline{T},\overline{\mathbb{Q}}),$
where
    $\overline{X}=X\times\Omega\times[-c,c]$ and $\overline{B}=\mathcal{B}\otimes\mathcal{A}\times\mathcal{B}([-c,c])$.
    $\overline{T}(x,\omega,t)=(\hat{T}^{\rho_{E}(x,t)}(x,\omega),t+f(x)+\cdots+f\circ T^{\rho_{E}(x,t)-1}(x))$.
    ¯=×$\overline{\mathbb{Q}}=\mathbb{Q}\times\lambda_{c}$.
It is easy to verify that $\overline{\mathbb{Q}}$ is $\overline{T}$-invariant. Note that the projection of $\overline{\mathbb{Q}}$ onto $E=X\times[-c,c]$ is $\theta_{E}$ and the disintegration of $\overline{\mathbb{Q}}$ with respect to $\theta_{E}$ is ¯$\overline{\mathbb{Q}}_{x,t}=\mathbb{Q}_{x}$ for $\theta_{E}$-a.e. $(x,t)\in E$.
Proof of Theorem 1.3.
We shall use the filling scheme [6]. For $g\in\mathbb{R}$ denote by $g^{+}=\max\{g,0\}$ and $g^{-}=\max\{-g,0\}$.
We may view both $f$ and $F$ as functions on $\overline{X}$ by taking $f(x,\omega,t)=f(x)$ and $F(x,\omega,t)=F(x,\omega)$. For $n\geq 1$ define
$G_{n}:=\max_{1\leq k\leq n}\overline{S}_{k}F,$
where we denote by
$\overline{S}_{k}F=F+F\circ\overline{T}+\cdots+F\circ\overline{T}^{k-1}.$
Then we have
$G_{n+1}^{+}-G_{n}^{+}\circ\overline{T}=F+G_{n+1}^{-}.$
Let $G=\lim_{n\to\infty}G_{n}$. Assume that ¯$\overline{\mathbb{Q}}(G<\infty)=1$. Then we may write
$G^{+}-G^{+}\circ\overline{T}=F+G^{-}.$
Let
$\kappa:=\inf\{n\geq 1:F\circ\overline{T}^{n-1}>0\}.$
Then, since for all $n\geq 1$, $G_{n}\circ\overline{T}^{\kappa-1}\geq F\circ\overline{T}^{\kappa-1}>0$, we get that $G\circ\overline{T}^{\kappa-1}\geq F\circ\overline{T}^{\kappa-1}>0$ hence $G^{+}\circ\overline{T}^{\kappa-1}=G\circ\overline{T}^{\kappa-1}$ and $G^{-}\circ\overline{T}^{\kappa-1}=0$. Denote by $H(x,\omega,t)=G^{\kappa(x,\omega,t)-1}(x,\omega,t)$ and $h(x,\omega,t)=F^{\kappa(x,\omega,t)-1}(x,\omega,t)$. Then we deduce that
$H-H\circ\overline{T}=h.$
For $u\in\mathbb{R}$ denote by
$\Phi_{u}(x):=\mathbb{E}_{\overline{\mathbb{Q}}_{x,t}}(e^{iuH(x,\cdot,t)})$
and
$\phi_{u}(x):=\mathbb{E}_{\overline{\mathbb{Q}}_{x,t}}(e^{iuh(x,\cdot,t)}).$
Then by (A) we have
$\Phi_{u}(x,t)=\phi_{u}(x,t)\cdot\Phi_{u}\circ T_{f}^{E}(x,t).$
Write $f(x,t)=-\log|\Phi_{u}(x,t)|$ and $g(x,t)=-\log|\phi_{u}(x,t)|$. Then we have that $f,g\geq 0$ and
$f-f\circ T_{f}^{E}=g.$
Take $N>0$ large enough such that
$E_{N}=E\cap\{f\leq N\}$
has positive $\theta_{E}$-mass. Let $(E_{N},\mathcal{B}_{N},T_{N},\theta_{N})$ denote the induced dynamical system by $E_{N}$. We get $f-f\circ T_{N}=g_{N}$, where
$g_{N}=g+g\circ T_{f}^{E}+\cdots+g\circ(T_{f}^{E})^{\rho_{E_{N}}(x,t)-1}.$
Now on $E_{N}$ we have that $0\leq f\leq N$, $0\leq f\circ T_{N}\leq N$ and $0\leq g_{N}\leq f\leq N$ are all bounded functions hence integrable, therefore by the $T_{N}$-invariance of $\mu_{N}$ and Birkhorff ergodic theorem,
$\mathbb{E}_{\theta_{N}}(g_{N}\,|\,\mathcal{I}_{N})(x,t)=0$
for $\theta_{N}$-a.e. $(x,t)\in E_{N}$, where $\mathcal{I}_{N}$ is the $T_{N}$-invariant $\sigma$-algebra of $E_{N}$. Since $g$ is non-negative function, $g_{N}$ is also nonnegative, we deduce that $g_{N}(x,t)=0$ and therefore $g(x,t)=0$ for $\theta_{N}$-a.e. $(x,t)\in E_{N}$. Since $N$ is arbitrary, we deduce that $g(x,t)=0$ or equivalently $|\phi_{u}(x,t)|=1$ for $\theta$-a.e. $(x,t)\in E$. This implies that for $\theta_{E}$-a.e. $(x,t)\in E$, $h(x,\cdot,t)=F^{\kappa(x,\cdot,t)-1}(x,\cdot,t)$ is $\mathbb{Q}_{x,t}$-a.s. a constant. But this is not possible under (A), the reason is the following: Recall the disjoint intervals $I_{x}$ and $J_{x}$ from (A). We may choose the intervals a bit smaller if necessary so that $a_{x}=\inf I_{x}\cup J_{x}>0$. Then we have $\mathbb{Q}_{x}(F(x,\cdot)\geq a_{x})>0$. Note that when $F(x,\omega)\geq a_{x}$, we have $\kappa(x,\omega,t)=1$ and $F^{\kappa(x,\omega,t)-1}(x,\omega,t)=F(x,\omega)$. Then by the fact that both $\mathbb{Q}_{x}(F(x,\cdot)\in I_{x})>0$ and $\mathbb{Q}_{x}(F(x,\cdot)\in J_{x})>0$, as well as $I_{x}$ and $J_{x}$ are disjoint, $F^{\kappa(x,\cdot,t)-1}(x,\cdot,t)$ can not be $\mathbb{Q}_{x,t}$-a.s. a constant.
This contradiction implies ¯$\overline{\mathbb{Q}}(G<\infty)<1$, or in other words, for $\theta_{E}$-positive $(x,t)\in E$, for $\mathbb{Q}_{x,t}$-positive $\omega\in\Omega$,
$\limsup_{n\to\infty}\overline{S}_{n}F(x,\omega,t)=+\infty.$
On the other hand, note that by construction, for $(x,t)\in E=X\times[-c,c]$, when $T_{f}^{E}(x,t)$ returns to $E$, we have
$t+f(x)+\cdots+f\circ T^{\rho_{E}(x,t)-1}(x)\in[-c,c].$
Thus for all $n\geq 1$,
$|\overline{S}_{n}f(x,\omega,t)|\leq 2c.$
Since $\theta_{E}=\mu\times\lambda_{c}$, we obtain that for $\mu$-positive $x\in X$, for $\mathbb{Q}_{x}$-positive $\omega$, there exists a subsequence $n_{k}=n_{k}(x,\omega)\to\infty$ such that $S_{n_{k}}f(x)\in[-2c,2c]$ for all $k\geq 1$ and $\limsup_{k\to\infty}S_{n_{k}}F(x,\omega)=+\infty$, therefore
$\limsup_{n\to\infty}S_{n}f(x)+S_{n}F(x,\omega)=+\infty.$