### v2

# A Chung-Fuchs type theorem for skew product dynamical systems

Abstract.
We prove a Chung-Fuchs type theorem for skew product dynamical systems such that for a measurable function on such a system, if its Birkhoff average converges to zero almost surely, and on typical fibres its Birkhoff sums have a non-trivial independent structure, then its associated generalised random walk oscillates, that is the supremum of the random walk equals to +$\infty$+ ∞ and the infimum equals to $\infty$- ∞.
1. Introduction
Let $(\Omega,\mathcal{A},\mathbb{P})$ be a probability space. Let $\{W_{n}\}_{n\geq 1}$ be a sequence of independent and identically distributed (i.i.d.) random variables in $\mathbb{R}$. Let
$S=\{S_{n}=W_{1}+\cdots+W_{n}\}_{n\geq 1}$
denote its associated random walk. We say that $S$ is recurrent if for all $c>0$,
$\mathbb{P}(|S_{n}|\leq c\text{ for infinitely many }n)=1.$
Regarding the recurrence of random walks we have the following well-known theorem of Chung and Fuchs [3]:
Theorem 1.1 (Chung and Fuchs 1951).
If $\mathbb{E}(|W_{1}|)<\infty$ and $\mathbb{E}(W_{1})=0$, then $S$ is recurrent. Furthermore, if $\mathbb{P}(W_{1}\neq 0)>0$, then $S$ oscillates, that is
(1.1) $\mathbb{P}\big{(}\liminf_{n\to\infty}S_{n}=-\infty\big{)}=\mathbb{P}\big{(}\limsup_{n\to\infty}S_{n}=+\infty\big{)}=1.$
In [3], the condition “$\mathbb{E}(|W_{1}|)<\infty$ and $\mathbb{E}(W_{1})=0$” is only to ensure that, by the strong law of large numbers,
(1.2) $\mathbb{P}\Big{(}\lim_{n\to\infty}\frac{S_{n}}{n}=0\Big{)}=1.$
So Chung and Fuchs’ theorem can be stated in a slightly more general way without $L^{1}$-integrable assumption that
(1.2)$\ \Rightarrow\ S\text{ is recurrent}.$
For (1.1), Chung and Fuchs showed that the set of recurrent points is a closed additive subgroup of $\mathbb{R}$, hence it is either empty, or the whole space $\mathbb{R}$, or a lattice $a\cdot\mathbb{Z}$ for some $a\in\mathbb{R}$. So if $S$ is recurrent, then the set of recurrent points is non-empty, thus it is either $\mathbb{R}$ or $a\cdot\mathbb{Z}$ for some $a\neq 0$ (since $\mathbb{P}(W_{1}\neq 0)>0$), and either case implies (1.1).
One may consider the Chung-Fuchs theorem for more general random walks: Let $(X,\mathcal{B},T,\mu)$ be a measure-preserving dynamical system (m.p.d.s) with $\mu(X)=1$. Let $f:X\to\mathbb{R}$ be a measurable function. Denote by
$S_{f}=\{S_{n}f=f+f\circ T+\cdots+f\circ T^{n-1}\}_{n\geq 1}$
its associated sequence of Birkhoff sums, or in other words, its associated generalised random walk. We say that $S_{f}$ is recurrent if for all $c>0$,
$\mu(|S_{n}f|\leq c\text{ for infinitely many }n)=1.$
The following extension of Chung-Fuchs theorem to generalised random walks is essentially due to Dekking [4] (see also [1] when assuming $f\in L^{1}(\mu)$).
Theorem 1.2 (Dekking 1982).
If $\mu$ is ergodic and
(1.3) $\mu\Big{(}\lim_{n\to\infty}\frac{S_{n}f}{n}=0\Big{)}=1,$
then $S_{f}$ is recurrent.
For generalised random walks, the statement that (1.3) plus $f\not\equiv 0$ implies (1.1) is no longer true. A simple counter example is the following: Let $X=\{x_{0},x_{1}\}$ be a space of two points and let the transformation $T$ be $T(x_{0})=x_{1}$, $T(x_{1})=x_{0}$ and let $\mu(\{x_{0}\})=\mu(\{x_{1}\})=\frac{1}{2}$. Take $f:X\to\mathbb{R}$ to be $f(x_{0})=1$, $f(x_{1})=-1$. Then $S_{n}f$ is periodic and (1.3) holds, but both liminf and limsup of $S_{n}f$ are bounded.
It is natural to ask for which settings between random walks and generalised random walks it holds that (1.3) plus $f\not\equiv 0$ implies (1.1). In [2] the authors studied the action of Mandelbrot cascades on ergodic measures, and in the estimation of the degeneracy of this action it was sufficient to show that
(1.4) $\mathbb{Q}\Big{(}\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty\Big{)}=1,$
where, in brief, the probability measure $\mathbb{Q}$ is a skew-product extension of the ergodic measure $\mu$ to $X\times\Omega$, $S_{n}f(\cdot)$ is a generalised random walk under $\mu$ and $S_{n}F(x,\cdot)$ is a random walk under $\mathbb{Q}_{x}$ for $\mu$-a.e. $x$, where the probability measure $\mathbb{Q}_{x}$ on $\Omega$ is the disintegration of $\mathbb{Q}$ w.r.t. $\mu$. In the sub-critical case when $S_{n}f+S_{n}F$ has a positive drift, that is $\mathbb{Q}$-a.s. $\lim_{n\to\infty}\frac{S_{n}f+S_{n}F}{n}=c>0$, it is easy to deduce (1.4). The critical case when $\mathbb{Q}(\lim_{n\to\infty}\frac{S_{n}f+S_{n}F}{n}=0)=1$ corresponds to (1.3) and it becomes more delicate to verify (1.4), which has led us to derive the following result (a more general and more abstract version of the result in [2]).
Main result
Recall that $(X,\mathcal{B},T,\mu)$ is a m.p.d.s with $\mu(X)=1$. Consider a skew-product measure-preserving dynamical system $(X\times\Omega,\hat{\mathcal{B}},\hat{T},\mathbb{Q})$ where
    ^$\hat{\mathcal{B}}=\mathcal{B}\otimes\mathcal{A}$.
    $\hat{T}(x,\omega)=(T(x),g_{x}(\omega))$ is a skew-product transformation, where for $x\in X$, $g_{x}:\Omega\to\Omega$ is $\mathcal{A}$-measurable and for $A\in\mathcal{A}$, $x\to g_{x}(A)$ is $\mathcal{B}$-measurable.
    $\Pi_{X}^{*}(\mathbb{Q})=\mu$, where $\Pi_{X}$ is the projection from $X\times\Omega$ onto $X$.
Consider the measurable partition
$\eta=\{\Pi_{X}^{-1}(x)=\{x\}\times\Omega:x\in X\}$
and let $\mathbb{Q}^{\eta}_{x,\omega}$ denote the conditional measure of $\mathbb{Q}$ w.r.t. $\eta$. For $\mathbb{Q}$-a.e. $(x,\omega)$, $\mathbb{Q}^{\eta}_{x,\omega}$ is a probability measure carried by $\{x\}\times\Omega$. Note that $\mathbb{Q}^{\eta}_{x,\omega}$ only depends on $x$. Hence, denoting $\Pi_{\Omega}$ the projection from $X\times\Omega$ onto $\Omega$, we may define a family of probability measures $\mathbb{Q}_{x}=\Pi_{\Omega}^{*}(\mathbb{Q}^{\eta}_{x,\omega})$ on $\Omega$ for $\mu$-a.e. $x\in X$ such that for any measurable function $F:X\times\Omega\to\mathbb{R}$,
$\int_{X\times\Omega}F(x,\omega)\,\mathbb{Q}(\mathrm{d}(x,\omega))=\int_{X}\int_{\Omega}F(x,\omega)\,\mathbb{Q}_{x}(\mathrm{d}\omega)\mu(\mathrm{d}x).$
The family of probability measures $\mathbb{Q}_{x}$ for $\mu$-a.e. $x\in X$ is also referred as the disintegration of $\mathbb{Q}$ with respect to $\mu$.
Let $F:X\times\Omega\to\mathbb{R}$ be a measurable function such that
  • (A)
    for $\mu$-a.e. $x\in X$,
    (i) $F(x,\cdot)$ is not $\mathbb{Q}_{x}$-a.s. a constant;
    (ii) the sequence $\{F(x,\cdot),F\circ\hat{T}(x,\cdot),F\circ\hat{T}^{2}(x,\cdot),\cdots\}$ is an independent sequence of random variables under $\mathbb{Q}_{x}$.
For $n\geq 1$ denote by
$\hat{S}_{n}F=F+F\circ\hat{T}+\cdots+F\circ\hat{T}^{n-1}.$
We have the following extension of Chung and Fuchs theorem.
Theorem 1.3.
Assume (A). If
$\mathbb{Q}\Big{(}\lim_{n\to\infty}\frac{\hat{S}_{n}F}{n}=0\Big{)}=1,$
then
$\mathbb{Q}\Big{(}\liminf_{n\to\infty}\hat{S}_{n}F=-\infty\Big{)}>0\text{ and }\mathbb{Q}\Big{(}\limsup_{n\to\infty}\hat{S}_{n}F=+\infty\Big{)}>0.$
Remark 1.4.
If we further assume that $\mathbb{Q}$ is $\hat{T}$-ergodic, then, since $\{\liminf_{n\to\infty}\hat{S}_{n}F=-\infty\}$ and $\{\limsup_{n\to\infty}\hat{S}_{n}F=+\infty\}$ are $\hat{T}$-invariant sets, Theorem 1.3 implies
$\mathbb{Q}\Big{(}\liminf_{n\to\infty}\hat{S}_{n}F=-\infty\Big{)}=\mathbb{Q}\Big{(}\limsup_{n\to\infty}\hat{S}_{n}F=+\infty\Big{)}=1.$
Remark 1.5.
Assumption (A) holds for a more general class of skew-product measures $\mathbb{Q}$ than the Peyrière measure studied in [2]. In particular we are not requiring that the law of $F(x,\cdot)$ under $\mathbb{Q}_{x}$ to be independent of $x$.
Remark 1.6.
We improved the proofs comparing to [2]. Firstly we used a result of Kesten to remove the method of using the stopping time that $F\circ\hat{T}^{n}(x,\cdot)$ firstly becomes positive, then we added an induced system argument to deduce that a measurable (not necessarily integrable) coboundary $u-u\circ T=v$ with $u,v\geq 0$ is equal to zero almost everywhere.
Example
We may consider the following canonical example: let $(X\times Y,\mathcal{B}_{X\times Y},\tilde{T},\Theta)$ be a skew product dynamical system of $(X,\mathcal{B},T,\mu)$, where the skew product $\tilde{T}(x,y)=(T(x),\varphi_{x}(y))$ is defined for a family of measurable functions $\varphi_{x}:Y\to Y$ and the $\tilde{T}$-invariant measure $\Theta$ satisfies $\Pi_{X}^{*}(\Theta)=\mu$. For $\mu$-a.e. $x\in X$ let $\Theta_{x}$ denote the disintegration of $\Theta$ w.r.t. $\mu$. We define the skew-product dynamical system $(X\times\Omega,\hat{\mathcal{B}},\hat{T},\mathbb{Q})$ as follows: Let $\Omega$ be the infinite product space of $Y$, that is,
$\Omega=\bigotimes_{\mathbb{N}}Y.$
The family of measurable functions $g_{x}:\Omega\to\Omega$ used to define the skew product $\hat{T}(x,\omega)=(T(x),g_{x}(\omega))$ is taken to be the left-shift operator $\sigma$, that is,
$g_{x}(\omega)=\sigma(\omega)=(y_{2},y_{3},\cdots)\text{ for }\omega=(y_{n})_{n\geq 1}\in\Omega.$
The $\hat{T}$-invariant measure $\mathbb{Q}$ on $X\times\Omega$ is defined as
$\mathbb{Q}(\mathrm{d}x,\mathrm{d}\omega)=\mu(\mathrm{d}x)\mathbb{Q}_{x}(\mathrm{d}\omega),$
where the fibre measures $\mathbb{Q}_{x}$ for $\mu$-a.e. $x\in X$ is the product measure
$\mathbb{Q}_{x}(\mathrm{d}\omega)=\prod_{k=1}^{\infty}\Theta_{T^{k-1}(x)}(\mathrm{d}y_{k})\text{ for }\omega=(y_{n})_{n\in\mathbb{N}}\in\Omega.$
Let $h:X\times Y\mapsto\mathbb{R}$ be a measurable function. Define the measurable function $F:X\times\Omega\to\mathbb{R}$ by
$F(x,\omega)=h(x,y_{1})\text{ for }x\in X\text{ and }\omega=(y_{n})_{n\in\mathbb{N}}\in\Omega.$
Then for $k\geq 1$ we have
$F\circ\hat{T}^{k-1}(x,\omega)=h(T^{k-1}(x),y_{k})\text{ for }x\in X\text{ and }\omega=(y_{n})_{n\in\mathbb{N}}\in\Omega.$
By the product structure it is straightforward that the sequence of random variables
$\{F(x,\cdot),F\circ\hat{T}(x,\cdot),F\circ\hat{T}^{2}(x,\cdot),\cdots\}$
is independent under $\mathbb{Q}_{x}$, therefore in this setting we have the following corollary from Theorem 1.3 (for simplicity we assume $h$ to be in $L^{1}$ and $\mathbb{Q}$ is ergodic).
Corollary 1.7.
Assume that
  • for $\mu$-a.e. $x\in X$, $h(x,\cdot)$ is not $\Theta_{x}$-a.s. a constant;
  • $h\in L^{1}(\Theta)$, $\int_{X\times Y}h\,\mathrm{d}\Theta=0$ and $\mathbb{Q}$ is $\hat{T}$-ergodic.
Then for $\mu$-a.e. $x\in X$ and for $\prod_{k=1}^{\infty}\Theta_{T^{k-1}(x)}$-a.e. $(y_{k})_{k\in\mathbb{N}}\in\bigotimes_{\mathbb{N}}Y$,
$\liminf_{n\to\infty}\sum_{k=1}^{n}h(T^{k-1}(x),y_{k})=-\infty\text{ and }\limsup_{n\to\infty}\sum_{k=1}^{n}h(T^{k-1}(x),y_{k})=+\infty.$
2. Proof of Theorem 1.3
We shall use the filling scheme, see [5] for example.
For $a\in\mathbb{R}$ denote by $a^{+}=\max\{a,0\}$ and $a^{-}=\max\{-a,0\}$.
For $n\geq 1$ define
$G_{n}:=\max_{1\leq k\leq n}\hat{S}_{k}F.$
We have
$G_{n+1}^{+}-G_{n+1}^{-}=G_{n+1}=F+G_{n}^{+}\circ\hat{T}.$
Let $G=\lim_{n\to\infty}G_{n}$. Then
(2.1) $G^{+}-G^{-}=F+G^{+}\circ\hat{T}.$
From the fact
$\mathbb{Q}\Big{(}\lim_{n\to\infty}\frac{\hat{S}_{n}F}{n}=0\Big{)}=1$
we deduce that, $\mathbb{Q}$-a.s., $\hat{S}_{n}F$ changes signs in the wide sense infinite many times, that is, $\hat{S}_{n}F>0$ or $\hat{S}_{n}F<0$ cannot hold for all $n$ large enough. This is due to a result of Kesten [6] that the sums of stationary sequences cannot grow slower than linearly, see [5, Section 5.(c)] for example. This implies that, $\mathbb{Q}$-a.s., $G\geq 0$ hence $G^{-}=0$, thus we may write from (2.1) that, $\mathbb{Q}$-a.s.,
(2.2) $G^{+}=F+G^{+}\circ\hat{T}.$
Assume that $\mathbb{Q}(G^{+}<\infty)=1$. This implies that for $\mu$-a.e. $x\in X$,
$\mathbb{Q}_{x}(G^{+}(x,\cdot)<\infty)=1.$
Fix $t\in\mathbb{R}\setminus\{0\}$. For $\mu$-a.e. $x\in X$ denote by
$\Phi_{t}(x):=\mathbb{E}_{\mathbb{Q}_{x}}(e^{itG^{+}(x,\cdot)})\text{ and }\phi_{t}(x):=\mathbb{E}_{\mathbb{Q}_{x}}(e^{itF(x,\cdot)}).$
Then by (2.2) and (A)(ii) we have, for $\mu$-a.e. $x\in X$,
$\Phi_{t}(x)=\phi_{t}(x)\times\Phi_{t}\circ T(x).$
Write $u(x)=-\log|\Phi_{t}(x)|$ and $v(x)=-\log|\phi_{t}(x)|$ with the convention that $-\log 0=\infty$. Then we have, $\mu$-a.s., $u,v\geq 0$ and
$u=v+u\circ T.$
Suppose that $\mu(u<\infty)>0$. Take $N>0$ large enough such that
$X_{N}=X\cap\{u\leq N\}$
has positive $\mu$-mass. Let $(X_{N},\mathcal{B}_{N},T_{N},\mu_{N})$ denote the induced dynamical system of $(X,\mathcal{B},T,\mu)$ by $X_{N}$, where $\mathcal{B}_{N}$ is $\mathcal{B}$ restricted to $X_{N}$, $\mu_{N}=\frac{1}{\mu(X_{N})}\mu|_{X_{N}}$ is the normalised $\mu$ restricted to $X_{N}$, and, denoting $\rho_{N}(x):=\inf\{n\geq 1:T^{n}(x)\in X_{N}\}$ the first visiting time to $X_{N}$, $T_{N}(x)=T^{\rho_{X_{N}}(x)}(x)$ is the induced transform from $X_{N}$ to $X_{N}$.
On $(X_{N},\mathcal{B}_{N},T_{N},\mu_{N})$ we obtain the co-boundary equation
$u-u\circ T_{N}=v_{N},$
where
$v_{N}(x)=v(x)+v\circ T(x)+\cdots+v\circ T^{\rho_{X_{N}}(x)-1}(x).$
On $X_{N}$ we have $0\leq u\leq N$, $0\leq u\circ T_{N}\leq N$ and $0\leq v_{N}\leq u\leq N$ are all bounded functions hence integrable. Thus by the $T_{N}$-invariance of $\mu_{N}$ and Birkhorff ergodic theorem we deduce that
$\mathbb{E}_{\mu_{N}}(v_{N}\,|\,\mathcal{I}_{N})(x)=0$
for $\mu_{N}$-a.e. $x\in X_{N}$, where $\mathcal{I}_{N}$ is the $T_{N}$-invariant $\sigma$-algebra of $X_{N}$. Since $v$ is non-negative, $v_{N}$ is also nonnegative, we deduce that $v_{N}(x)=0$ and therefore $v(x)=0$ for $\mu_{N}$-a.e. $x\in X_{N}$. Since $N$ is arbitrary, we deduce that $v(x)=0$ or equivalently $|\phi_{t}(x)|=1$ for $\mu$-a.e. $x\in X\cap\{u<\infty\}$. Note that $|\phi_{t}(x)|=1$ implies $F(x,\cdot)$ is $\mathbb{Q}_{x}$-a.s. a constant, which is contradict to (A)(i). Therefore $\mu(u<\infty)=0$, or in other words,
(2.3) $\mu(x\in X:\Phi_{t}(x)=0)=1.$
We may take a countable sequence $t_{n}$ tending to 00. Then (2.3) implies that
$\mu(x\in X:\Phi_{t_{n}}(x)=0\text{ for all }n\geq 1)=1.$
But this is not possible since for $\mu$-a.e. $x\in X$, $\mathbb{Q}_{x}(G^{+}(x,\cdot)<\infty)=1$, that is, $G^{+}(x,\cdot)$ is a proper random variable under $\mathbb{Q}_{x}$, thus $\mathbb{E}_{\mathbb{Q}_{x}}(e^{itG^{+}(x,\cdot)})$ is continuous in $t$, hence non-vanishing around 00.
We finally deduce that $\mathbb{Q}(G^{+}<\infty)<1$, or in other words,
$\mathbb{Q}\Big{(}\limsup_{n\to\infty}\hat{S}_{n}F=+\infty\Big{)}>0.$
Since same arguments apply to $-F$, we also get
$\mathbb{Q}\Big{(}\liminf_{n\to\infty}\hat{S}_{n}F=-\infty\Big{)}>0.$